# Dreamweaver: Learning Compositional World Representations from Pixels

### v2

@@ Abstract
Humans have an innate ability to decompose their perceptions of the world into objects and their attributes, such as colors, shapes, and movement patterns. This cognitive process enables us to imagine novel futures by recombining familiar concepts. However, replicating this ability in artificial intelligence systems has proven challenging, particularly when it comes to modeling videos into compositional concepts and generating unseen, recomposed futures without relying on auxiliary data, such as text, masks, or bounding boxes. In this paper, we propose Dreamweaver, a neural architecture designed to discover hierarchical and compositional representations from raw videos and generate compositional future simulations. Our approach leverages a novel Recurrent Block-Slot Unit (RBSU) to decompose videos into their constituent objects and attributes. In addition, Dreamweaver uses a multi-future-frame prediction objective to capture disentangled representations for dynamic concepts more effectively as well as static concepts. In experiments, we demonstrate our model outperforms current state-of-the-art baselines for world modeling when evaluated under the DCI framework across multiple datasets. Furthermore, we show how the modularized concept representations of our model enable compositional imagination, allowing the generation of novel videos by recombining attributes from different objects.
@@ 1 Introduction
We achieve this goal by introducing a neural architecture named Dreamweaver. Dreamweaver encodes a temporal context window of $T$ past images using a novel Recurrent Block-Slot Unit (RBSU). The RBSU represents its state as a set of slot states, each updated independently. This is then passed through a block-slot bottleneck, mapping the potentially entangled monolithic slot state into a composition of independent blocks. Since a block vector can only obtain its value through attention to the prototype concept library, this process can also be seen as implementing an attraction process at each step. At the end of the temporal encoding, Dreamweaver predicts the observation at a future time step from the latest block-slot representation. We found that this predictive reconstruction objective is indeed crucial in making the dynamic concept abstractions emerge. In experiments, Dreamweaver outperforms state-of-the-art object-centric methods under the DCI framework across several datasets. Moreover, we demonstrate how RBSU’s modularized concept representations enable compositional imagination, generating novel videos that combine attributes from different objects.
@@ 2.2 Independent Block Attractor Dynamics
To maintain modularized representations from GRU and MLP, we utilize a separate library of prototypes ${\bf C}_{m}\in\mathbb{R}^{N_{\text{prototypes}}\times D}$ for each $m$. Notably, each library is initialized with $N_{\text{prototypes}}$ learnable vectors, which are subsequently learned through backpropagation to capture emergent semantic concepts.
@@ 3 Dreamweaver: Compositional World Model via Predictive Imagination
In this section, we propose and describe a novel compositional world model called Dreamweaver using the proposed RBSU. Broadly, Dreamweaver takes a video ${\bf x}_{1},\ldots,{\bf x}_{T}$ as input, constructs the world state in terms of composable tokens, and predicts $K$ future video frames $\hat{{\bf x}}_{T+1},\ldots,\hat{{\bf x}}_{T+K}$:
@@
where the encoder $f_{\phi}$ encodes the video ${\bf x}_{1},\ldots,{\bf x}_{T}$ into slots ${\bf S}_{T}$ leveraging RBSU. Given the slots ${\bf S}_{T}$ and a step indicator $k$, where $k$ is the relative temporal index of the target frame, the decoder $g_{\theta}$ predicts the target frame $\hat{{\bf x}}_{T+k}$ via an autoregressive image transformer (Singh et al., 2022a; b; 2023).
@@ C.3 Baselines
  • STEVE (Singh et al., 2022b): https://github.com/singhgautam/steve
  • SysBinder (Singh et al., 2023): https://github.com/singhgautam/sysbinder
  • RSSM (Hafner et al., 2020): https://github.com/jurgisp/pydreamer

### v3

# Dreamweaver: Learning Compositional World Models from Pixels

@@ Abstract
Humans have an innate ability to decompose their perceptions of the world into objects and their attributes, such as colors, shapes, and movement patterns. This cognitive process enables us to imagine novel futures by recombining familiar concepts. However, replicating this ability in artificial intelligence systems has proven challenging, particularly when it comes to modeling videos into compositional concepts and generating unseen, recomposed futures without relying on auxiliary data, such as text, masks, or bounding boxes. In this paper, we propose Dreamweaver, a neural architecture designed to discover hierarchical and compositional representations from raw videos and generate compositional future simulations. Our approach leverages a novel Recurrent Block-Slot Unit (RBSU) to decompose videos into their constituent objects and attributes. In addition, Dreamweaver uses a multi-future-frame prediction objective to capture disentangled representations for dynamic concepts more effectively as well as static concepts. In experiments, we demonstrate our model outperforms current state-of-the-art baselines for world modeling when evaluated under the DCI framework across multiple datasets. Furthermore, we show how the modularized concept representations of our model enable compositional imagination, allowing the generation of novel videos by recombining attributes from previously seen objects. cun-bjy.github.io/dreamweaver-website
@@ 1 Introduction
We achieve this goal by introducing a neural architecture named Dreamweaver. Dreamweaver encodes a temporal context window of $T$ past images using a novel Recurrent Block-Slot Unit (RBSU). The RBSU represents its state as a set of slot states, each updated independently. This is then passed through a block-slot bottleneck, mapping the potentially entangled monolithic slot state into a composition of independent blocks. Since a block vector can only obtain its value through attention to the prototype concept library, this process can also be seen as implementing an attraction process at each step. At the end of the temporal encoding, Dreamweaver predicts the observation at a future time step from the latest block-slot representation. We found that this predictive reconstruction objective is indeed crucial in making the dynamic concept abstractions emerge. In experiments, Dreamweaver outperforms state-of-the-art object-centric methods under the DCI framework across several datasets. Moreover, we demonstrate how RBSU’s modularized concept representations enable compositional imagination, generating novel videos that combine attributes from previously seen objects.
@@ 2.2 Independent Block Attractor Dynamics
To maintain modularized representations from GRU and MLP, we utilize a separate library of prototypes ${\bf C}_{m}\in\mathbb{R}^{N_{\text{prototypes}}\times D}$ for each $m$. Notably, each library is initialized with $N_{\text{prototypes}}$ learnable vectors, which start with random values and are subsequently learned through backpropagation to capture emergent semantic concepts.
@@ 3 Dreamweaver: Compositional World Model via Predictive Imagination
In this section, we propose and describe a novel compositional world model called Dreamweaver using the proposed RBSU. Broadly, Dreamweaver takes video frames ${\bf x}_{1},\ldots,{\bf x}_{T}$ as input, constructs the world state in terms of composable tokens, and predicts $K$ future video frames $\hat{{\bf x}}_{T+1},\ldots,\hat{{\bf x}}_{T+K}$:
@@
where the encoder $f_{\phi}$ encodes the video ${\bf x}_{1},\ldots,{\bf x}_{T}$ into slots ${\bf S}_{T}$ leveraging RBSU. Given the slots ${\bf S}_{T}$ and a step indicator $k$, where $k$ is the relative temporal index of the target frame, the decoder $g_{\theta}$ predicts the target frame $\hat{{\bf x}}_{T+k}$ via an autoregressive image transformer (Chen et al., 2020; Singh et al., 2022a; b; 2023).